Find the maximum value of y=-x^2+6x.
Answer by venugopalramana(3286)

Find the maximum value of y=-x^2+6x. =-{X^2-2*X*3+3^2}+9=9-(X-3)^2
HENCE AT X=3 ,WE GET MAXIMIUM VALUE FOR Y =9